in a closed container, gas with an initial volume of 0.125L is heated from 32 degrees Celsius to 96 degrees Celsius at 1atm what is the final volume... please explain why
Answer #1

Given:
Vi = 0.125 L
Ti = 32.0 oC
= (32.0+273) K
= 305 K
Tf = 96.0 oC
= (96.0+273) K
= 369 K

use:
Vi/Ti = Vf/Tf
0.125 L / 305.0 K = Vf / 369.0 K
Vf = 0.1512 L
Answer: 0.151 L